i'm confused on this problem:
Judy needs a mean(average) socre of 86 on four tests to earn a midterm grade of B. If the the mean of her first three test was 83, what is the lowest score ona 100-point scale that she can receive on the fourth test to have a midterm grade of B? **Let 't' represent the unknown test grade

Answer by venugopalramana(3286)

Judy needs a mean(average) socre of 86 on four tests to earn a midterm grade of B. If the the mean of her first three test was 83, what is the lowest score ona 100-point scale that she can receive on the fourth test to have a midterm grade of B? **Let 't' represent the unknown test grade
TOTAL TESTS =4
AVERAGE NEEDED TO GET B GRADE = 86
SO TOTAL MARKS NEEDED IN 4 TESTS =AVERAGE *NUMBER OF TESTS=86*4=344
NUMBER OF TESTS ATTENDED BY JUDY =3
AVERSGE GOT IN 3 TESTS =83
SO TOTAL MARKS GOT IN 3 TESTS =AVERAGE * NUMBER OF TESTS=83*3= 249
SO MARKS TO BE GOT IN 4 TH. TEST =344-249= 95